# QN-Mixer: A Quasi-Newton MLP-Mixer Model for Sparse-View CT Reconstruction

### v1

@@ 3 Reproducibility
All our experiments are fully reproducible. While the complete algorithm is already provided in the main paper (see Algorithm 2), we additionally present a PyTorch pseudo-code for enhanced reproducibility in Sec. 3.1. We furnish comprehensive references to all external libraries used in Sec. 3.2. Detailed information regarding the initialization of our model can be found in Sec. 3.3. The precise parameters of our regularizer, Incept-Mixer architecture, are available in Sec. 3.4. We outline the exact data splits utilized across the paper for the AAPM dataset in Sec. 3.5. Lastly, to facilitate the reproduction of our out-of-distribution (OOD) protocol, we provide the pseudo-code in Sec. 3.6.
3.1 QN-Mixer pseudo-code
@@
3.2 External libraries used
@@
3.3 QN-Mixer’s parameters initialization
@@
3.4 Incept-Mixer’s architecture
@@ 17.6 $96\times 256\times 256$
3.5 AAPM dataset splits
@@
3.6 Robustness eval. protocol for OOD scenarios
@@
In the main paper, we propose a novel protocol specifically crafted for evaluating the effectiveness of reconstruction methods when handling abnormal data lying outside the distribution of the training dataset. Assessing the model’s capability to reconstruct abnormal data holds significant relevance in clinical applications, where patient data may exhibit deviations from the characteristics present in the training data. We strongly advocate for future research endeavors to embrace and employ this protocol as a standard for evaluating the robustness of reconstruction methods. To facilitate the seamless integration of this protocol, we furnish the function’s pseudo-code in Algorithm 4.
4 Limitations

### v3

@@ Appendix D Reproducibility
All our experiments are fully reproducible. While the complete algorithm is already provided in the main paper (see Algorithm 2), we additionally present a PyTorch pseudo-code for enhanced reproducibility in Sec. D.1. We furnish comprehensive references to all external libraries used in Sec. D.2. Detailed information regarding the initialization of our model can be found in Sec. D.3. The precise parameters of our regularizer, Incept-Mixer architecture, are available in Sec. D.4. We outline the exact data splits utilized across the paper for the AAPM dataset in Sec. D.5. Lastly, to facilitate the reproduction of our out-of-distribution (OOD) protocol, we provide the pseudo-code in Sec. D.6.
D.1 QN-Mixer pseudo-code
@@
D.2 External libraries used
@@
D.3 QN-Mixer’s parameters initialization
@@
D.4 Incept-Mixer’s architecture
@@ 17.6 $96\times 256\times 256$
D.5 AAPM dataset splits
@@
D.6 Robustness eval. protocol for OOD scenarios
@@
In medical imaging, it’s crucial to develop methods that generalize to scans with lesions or anomalies, and assessing the model’s capability to reconstruct abnormal data holds significant relevance, as test patient data may deviate from the training data in clinical applications. To this end, we design a simple protocol specifically crafted for evaluating the effectiveness of methods when handling abnormal data. In this case, a white circle mimicking an out-of-distribution texture, which was never seen during training, is forged into CT images with noise-free sinograms. The pseudo-code to realize this is provided in Algorithm 4. Then, performance can be evaluated on the entire image as shown in Tab. 3, or on a cropped region within the circle as detailed in Tab. 8.
We strongly advocate for future research endeavors to embrace and employ this protocol as a standard for evaluating the robustness of reconstruction methods.
Appendix E Limitations